1. As the ball drops, it has an equal chance of making a left turn or right turn at each peg.

SCOREBOARD INNING RUNS OUTS

A. What is the probability of a home run? 

P(home run)=

B. What is the probability of an out?

P(out)=

C. What is the probability of a hit (a single, double, triple, or home run)? 

P(hit)=

D. How are the answers to parts B and C related?
Answer #1

Answer

total number of possible outcome = 3 out + 1 home run + 1 single + 1 double +1 triple = 7

(A) favourable outcome for home run is only 1 and total outcome is 7

so, probability = (favourable outcome)/(total outcome)

P(home run) = 1/7

(B) favourable outcome for an out is 3 and total outcome is 7

so, probability = (favourable outcome)/(total outcome)

P(out) = 3/7

(C) favourable outcome for home run, single, double or triple is 4 ( or 1+1+1+1) and total outcome is 7

so, probability = (favourable outcome)/(total outcome)

P(hit) = 4/7

(D) answer to parts B and C are complement of each other or we can say that both answers make the complete area, i.e. 1

complement means opposite of each other